OpenStudy (anonymous):

how would you solve 3x+3+2x=4x? I have an idea, I want to see if I did it right

OpenStudy (anonymous):

Well how did you do it?

OpenStudy (anonymous):

3x+3-3+2x=4x-3 3x+2x=x

OpenStudy (anonymous):



OpenStudy (anonymous):

@corrina.howerton you cant add or subtract terms that are not alike. For example you cant subtract 3x - 3 = 0 but you can subtract 3x - 3x = 0 or 3 - 3 = 0

OpenStudy (anonymous):

I got -3 = x as well.

OpenStudy (anonymous):

I think I am finally understanding what I am doing on this stuff, at least I hope so I can get it done. Thank you both for your help

OpenStudy (anonymous):